OpenStudy (anonymous):

how do i find tan(theta symbol)=1 on the unit circle?

OpenStudy (amistre64):

find all angles that have a slope of 1; tan = m = y/x = slope

OpenStudy (anonymous):

oh how do i find these angles without randomly dividing the coordinates?

OpenStudy (amistre64):

trig is mostly memorizing stuff; and it helps to have memorized that y=x is a line that has a slope of 1

OpenStudy (amistre64):

and an angle of 45 degrees

OpenStudy (anonymous):

oo ya (√2/2,√2/2)

OpenStudy (amistre64):

since tan(t) = 1; slope = 1 ... thats it lol ;)

OpenStudy (amistre64):

2 angles can produce a slope of one tho .... 45 and 45+180

OpenStudy (anonymous):

wat u mean +180?

OpenStudy (anonymous):

y = sin t x= cos t y/x =tan t

OpenStudy (anonymous):

yes my answer requires 2 solutions though

OpenStudy (anonymous):

Tangent is positive in 3rd quadrant.

OpenStudy (anonymous):

oh so 5pi/4 is second one right?

OpenStudy (anonymous):

ie y= -sint x= - cos t y/x = tan t

OpenStudy (anonymous):

Yes, that's right.

OpenStudy (anonymous):

now what if for example cot(theta)=-√3 im no longer looking for when y=x do u just try random ones?